OpenStudy (misty1212):

if a chicken and a half can lay an egg and a half in a day and a half, how long will it take one chicken to lay two dozen eggs?

OpenStudy (alexandervonhumboldt2):

she will never lay them lol

OpenStudy (alexandervonhumboldt2):

jk

OpenStudy (anonymous):

wait how can a chicken lay half an egg

OpenStudy (anonymous):

oh nvm

OpenStudy (alexandervonhumboldt2):

it is impossible, that why i answered that she will never lay them

OpenStudy (alexandervonhumboldt2):

@ganeshie8

OpenStudy (anonymous):

o...k

OpenStudy (alexandervonhumboldt2):

but she can lie 1 egg

OpenStudy (anonymous):

lie

OpenStudy (alexandervonhumboldt2):

maybe round 1.5chiken to 1 chiken

OpenStudy (alexandervonhumboldt2):

and 1.5 egg to 1 egg

OpenStudy (anonymous):

yea im done questioning this

OpenStudy (alexandervonhumboldt2):

this problem is same as `there are 5 apples on 1 oak and 3 apples on other oak. how many apples there are?`

OpenStudy (anonymous):

that aint even logical question

OpenStudy (anonymous):

Actually this question kind of confused me. Is it 16?

OpenStudy (owlcoffee):

Let's call "chickens" for \(\delta\), eggs for \(\beta\) and days as \(\Omega\) it should imply, by the gven information that the relationship is: \[\frac{ 3 }{ 2 } \delta = \frac{ 3 }{ 2 } \beta = \frac{ 3 }{ 2 } \Omega\] And cancelling things out should yield: \[\delta = \beta = \Omega\] Which implies that they have a proportional relationship and the relationship we are looking for is 1:24 between shickens and eggs, that would allow us to calculate how many chicken and eggs there will be in some timeframe in such a way that every chicken would've laid 24 eggs. Those are my first thoughts on this.

OpenStudy (misty1212):

what the monkey is a \(\Omega\)?

OpenStudy (owlcoffee):

a horse-shoe. ... Okay, it's the greek letter "Omega".

OpenStudy (misty1212):

yeah i got the watch!

OpenStudy (misty1212):

gtg, i'll ask again later

OpenStudy (owlcoffee):

Have a nice day.

OpenStudy (anonymous):

wtf that was not even a logical question

Directrix (directrix):

Chicken and egg problem discussion: http://mathforum.org/library/drmath/view/58675.html @misty1212